OpenStudy (anonymous):

Which of the following equations represents an inverse variation? y equals x divided by two y equals two times x y equals two divided by x y equals x squared

OpenStudy (anonymous):

@BrAnDoN_AAS @cool41

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ y=\frac{k}{x} }\) is an inverse variation.

OpenStudy (solomonzelman):

k is a constant.

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ y=\frac{x}{2} }\) is same as \(\LARGE\color{black}{ y=\frac{1}{2} x }\) so it is not inverse. \(\LARGE\color{blue}{ y=2x }\) is obviously not an inverse variation, it is direct. \(\LARGE\color{red}{ y=\frac{2}{x} }\) is more like it. \(\LARGE\color{magenta}{ y=x^2 }\) is a parabola.

OpenStudy (anonymous):

y equals 2 divided by x

OpenStudy (solomonzelman):

So...

OpenStudy (anonymous):

@elliecan you have de answer

OpenStudy (solomonzelman):

y=2/x is correct. It is an inverse variation.

OpenStudy (anonymous):

but thats wrong

OpenStudy (anonymous):

@cool41

OpenStudy (anonymous):

i am pretty sure it is y =x/2. letme know if it is right

OpenStudy (anonymous):

ya thats what I meant first time but got it mixed up

OpenStudy (anonymous):

XD